OpenStudy (anonymous):

I need help on last question. Solve 2(x + 1) + 4 = 8. 1 6.5 4 3

OpenStudy (texaschic101):

2(x + 1) + 4 = 8 --- subtract 4 from both sides 2(x + 1) = 4 --- divide both sides by 2 x + 1 = 2 -- subtract 1 from both sides x = 1 check.. 2(x + 1) + 4 = 8 2(1 + 1) + 4 = 8 2(2) + 4 = 8 4 + 4 = 8 8 = 8 (correct) So x = 1 any questions ?

OpenStudy (anonymous):

No none at all you explained it perfectly thank you so much! :)

OpenStudy (texaschic101):

sure thing :)